OpenStudy (anonymous):

What is the 23rd term of the arithmetic sequence where a1 = 7 and a5 = 21?

OpenStudy (anonymous):

the equation for an arithmetic sequence is: an = a1 + (n – 1)d

OpenStudy (anonymous):

First we'll have to solve for d, since we don't know what the distance each term is added/subtracted. This case added.

OpenStudy (anonymous):

so, a(5)=7+(5-1)d where a(5) = 21 21=7+(5-1)d just solve for d and then use the arithmetic sequence equation to find the 23rd term :)

OpenStudy (anonymous):

thank you